# Flexible inference of evolutionary accumulation dynamics using uncertain observational data

### v2

@@ 4 Discussion
Some other aspects of our approach can be further generalised in future. The simulated annealing optimisation is a random process, so evaluating the results for different random seeds will enhance robustness at a cost of additional computational time. We assume (commonly with other evolutionary accumulation approaches) that features are accumulated irreversibly, a simplification that does not necessarily hold true in the actual underlying biological process [Renz et al., 2024]. For implementing this reversibility [Johnston and Diaz-Uriarte, 2024] in HyperLAU, further work is needed, but we think that our linear algebra approach could also be a reasonable starting point, for example by allowing also upper-diagonal entries in the transition matrix to be positive.
@@ Acknowledgments
This work was supported by the Trond Mohn Foundation [project HyperEvol under grant agreement No.
TMS2021TMT09 to IGJ], through the Centre for Antimicrobial Resistance in Western Norway (CAMRIA) [TMS2020TMT11].
@@
what means, that the assumption that the ancestor was sampled at $t=1$ cannot be true. The fact that both $c^{a}\star P_{.,2}$ and $c^{a}\star P_{.,3}$ are the null vectors indicate that the probability of the ancestor data point originate from step $t=2$ or $t=3$ is zero and we therefore don’t need to consider these option any further. Taking together the positive entries in all $c^{d}\star Q_{.,i}$ vectors and normalize them, we obtain the following likelihood of seeing the first data point given transition matrix $A$:
@@
Figure S1: Progression of the best log-likelihood found until then of the first toy example under the different models during the SA process. The iteration number on the x-axes indicates the iteration number in the optimization under model $F$.
Figure S2: Progression of the best log-likelihood of the second toy example the SA process. The iteration number on the x-axes indicates the iteration number in the optimization. The different colours of the trajectories indicate in which feature the uncertainty markers were inserted in the corresponding simulation. (top) Model $F$; (bottom) model 2.
Figure S3: The learned pathways for the second dataset used in section 3.1 under model $F$, but with approximately 50% randomly inserted uncertainty markers in A: feature 2, B: feature 3, C: feature 4, D: feature 5, E: feature 6. The strength of the flux is indicated by the edge-width and the CV is marked by the colour of the edges.
Figure S4: The learned pathways for the second dataset used in section 3.1 under model 2, but with approximately 40% randomly inserted uncertainty markers in A: feature 1, B: feature 2, C: feature 4, D: feature 5, E: feature 6. The strength of the flux is indicated by the edge-width and the CV is marked by the colour of the edges.

### v3

@@ 4 Discussion
Some other aspects of our approach can be further generalised in future. The simulated annealing optimisation is a random process, so evaluating the results for different random seeds will enhance robustness at a cost of additional computational time. We assume (commonly with other evolutionary accumulation approaches) that features are accumulated irreversibly, a simplification that does not necessarily hold true in the actual underlying biological process [Renz et al., 2025]. For implementing this reversibility [Johnston and Diaz-Uriarte, 2024] in HyperLAU, further work is needed, but we think that our linear algebra approach could also be a reasonable starting point, for example by allowing also upper-diagonal entries in the transition matrix to be positive.
@@ Acknowledgments
This work was supported by the Trond Mohn Foundation [project HyperEvol under grant agreement No. TMS2021TMT09 to IGJ], through the Centre for Antimicrobial Resistance in Western Norway (CAMRIA) [TMS2020TMT11].
@@
which means that the assumption that the ancestor was sampled at $t=1$ cannot be true. The fact that both $c^{a}\star P_{.,2}$ and $c^{a}\star P_{.,3}$ are the null vectors indicate that the probability of the ancestor data point originate from step $t=2$ or $t=3$ is zero and we therefore don’t need to consider these option any further.
Taking together the positive entries in all $c^{d}\star Q_{.,i}$ vectors and normalize them, we obtain the following likelihood of seeing the first data point given transition matrix $A$:
@@
Figure S1: Progression of the best log-likelihood of the first toy example during the SA process. The iteration number on the x-axes indicates the iteration number in the optimization under model $F$.
Figure S2: Progression of the best log-likelihood of the second toy example during the SA process. The iteration number on the x-axes indicates the iteration number in the optimization. The different colours of the trajectories indicate in which feature the uncertainty markers were inserted in the corresponding simulation. (top) Model $F$; (bottom) model 2.
Figure S3: Inferred pathways for the second dataset used in section 3.1 under model $F$, but with approximately 50% randomly inserted uncertainty markers in A: feature 2, B: feature 3, C: feature 4, D: feature 5, E: feature 6. The strength of the flux is indicated by the edge-width and the CV is marked by the colour of the edges.
Figure S4: Inferred pathways for the second dataset used in section 3.1 under model 2, but with approximately 40% randomly inserted uncertainty markers in A: feature 1, B: feature 2, C: feature 4, D: feature 5, E: feature 6. The strength of the flux is indicated by the edge-width and the CV is marked by the colour of the edges.